OpenStudy (anonymous):

compute the limit lim(8x-x^2-16)^1/2 as x goes to 4 my anserw was -32 is that right?

OpenStudy (anonymous):

substitute x by 4 in the expression: (8*4 -4^2-16)^1/2=(32-32)^1/2=0

OpenStudy (anonymous):

so 0 would be the anserw?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you